A departmental store has a single cashier. During the rush hours, customers arrive at the rate of 20 customers per hour. The average number of customers that can be processed by the cashier is 24 per hour.Calculate

a) Probability that the cashier is idle?b)Average number of customers in the queuing system?




1
Expert's answer
2021-03-25T06:32:22-0400

Mean arrival rate "a=15" customers per hour

Mean service rate "s=18" customers per hour.


a) "P_o=1-\frac{a} {s} =1-\frac{15}{18}=\frac {1} {6}"


b) "N=\frac {a} {s-a} =\frac {15}{3}=5"